# GFT: Graph Feature Tuning for Efficient Point Cloud Analysis

### v1

@@ 4.2 Classification
For fair and robust comparison, we reproduce the results from IDPT [38] and DAPT [44] under the same experimental setting as ours. In Table 1, we report the average overall accuracy (OA%) of models with five consecutive seeds within $[0,4]$. We design the task head to take inputs from CLS token $T_{cls}$ and pooled prompt and patch tokens from the last layer of the transformer. Two datasets are used for evaluations: ScanObjectNN for real-world object classification and ModelNet-40 for synthetic object classification.
ScanObjectNN [27] is a real-world point cloud dataset containing scanned objects from indoor scenes. It consists of approximately 15K object scans with background clutter and occlusions, categorized into 15 object types. We use three dataset splits: OBJ_BG, OBJ_ONLY, and PB_T50_RS. For preprocessing, we sample 2K points per instance and apply a simple rotation, inspired by ACT [5].
ModelNet-40 [34] consists of 12,311 CAD models of synthetic objects across 40 different categories. The point clouds in these models are clean, noise-free, and uniformly distributed with independent and identical distributions. Following the standard approach, we split the dataset into 9,843 instances for training and 2,468 for testing. During training, we sample 1K points from the distribution and apply random scaling and translation.
@@ 4.3 Segmentation
For segmentation, we design a transformer decoder with skip connections from the 3rd, 6th, 9th, and 11th layers of the pretrained model. Architecture used for the decoder of the segmentation is provide in the supplementary (Section 10). In line with the baseline papers IDPT [38] and DAPT [44], we report the best performance across ten consecutive seeds, $[0,9]$, as reproducing the baselines to compute mean metrics was computationally demanding. The evaluation dataset for the segmentation task is ShapeNetPart.
ShapeNetPart [36] is a part segmentation dataset containing 50 different parts across 15 object categories. Following the standard benchmarking protocol [44, 38], we divide the dataset into training and test splits, consisting of 14K and 2.9K scans, respectively. Each data instance is represented by 2K sampled points.
@@ 5 Results
The baseline foundational models favor clean and uniform point cloud data, similar to the pretraining dataset ShapeNet [2]. However, real-world point clouds have a different distribution due to noise, missing data, and non-uniform point density. Thus, we benchmark with ScanObjectNN [27].
As shown in Table 1, our method achieves an average performance gain of 4.12% over the FFT baseline of Point-BERT and 2.47% over Point-MAE. However, compared to the ACT baseline, there is an average performance drop of 0.5%. This could be attributed to ACT’s multi-stage pretraining [5], which, while effective, is computationally inefficient.
@@
ModelNet40 [34] contains synthetic models of objects. As illustrated in Table 1, our method does not exceed the FFT baselines of the pretrained backbones but remains competitive, with only a 0.5% average accuracy deficit. This finding of FFT can be attributed to the high similarity between the pretraining and fine-tuning datasets, as both consist of synthetic CAD models with a clean and uniform point distribution. Notably, our method outperforms other PEFT approaches across all three pretrained backbones.
@@
DGCNN-OcCo [29] 90.6 $\pm$ 2.8 92.5 $\pm$ 1.9 82.9 $\pm$ 1.3 86.5 $\pm$ 2.2
Transformer-OcCo [37] 94.0 $\pm$ 3.6 95.9 $\pm$ 2.3 89.4 $\pm$ 5.1 92.4 $\pm$ 4.6
MaskPoint [17] 95.0 $\pm$ 3.7 97.2 $\pm$ 1.7 91.4 $\pm$ 4.0 93.4 $\pm$ 3.5
EPCL[13] 95.1 $\pm$ 2.7 97.3 $\pm$ 1.6 91.1 $\pm$ 4.2 93.5 $\pm$ 3.8
Point-M2AE [39] 96.8 $\pm$ 1.8 98.3 $\pm$ 1.4 92.3 $\pm$ 4.5 95.0 $\pm$ 3.0
Point-MAE [19] 96.3 $\pm$ 2.5 97.8 $\pm$ 1.8 92.6 $\pm$ 4.1 95.0 $\pm$ 3.0
Point-BERT [37] 94.6 $\pm$ 3.1 96.3 $\pm$ 2.7 91.0 $\pm$ 5.4 92.7 $\pm$ 5.1
ACT [5] 96.8 $\pm$ 2.3 98.0 $\pm$ 1.4 93.3 $\pm$ 4.0 95.6 $\pm$ 2.8
@@
IDPT [38] 97.3 $\pm$ 2.1 97.9 $\pm$ 1.1 92.8 $\pm$ 4.1 95.4 $\pm$ 2.9
DAPT [44] 96.8 $\pm$ 1.8 98.0 $\pm$ 1.0 93.0 $\pm$ 3.5 95.5 $\pm$ 3.2
@@
IDPT [38] 96.0 $\pm$ 1.7 97.2 $\pm$ 2.6 91.9 $\pm$ 4.4 93.6 $\pm$ 3.5
DAPT [44] 95.8 $\pm$ 2.1 97.3 $\pm$ 1.3 92.2 $\pm$ 4.3 94.2 $\pm$ 3.4
@@
IDPT [38] 96.4 $\pm$ 2.5 98.3 $\pm$ 1.6 92.5 $\pm$ 4.5 95.6 $\pm$ 2.9
DAPT [44] 96.0 $\pm$ 3.0 97.4 $\pm$ 2.2 90.5 $\pm$ 4.9 94.2 $\pm$ 3.6
@@
Table 2: Few-shot learning on ModelNet40 [34]. We report the average overall accuracy (%) with the standard deviation (%) of 10 independent experiments.
Following previous works [44, 38, 5, 19, 37], we evaluate the pretrained model on N-way-K-shot few-shot classification. In this setting, $N\in\{5,10\}$ represents the number of object categories, while $K\in\{10,20\}$ denotes the number of training instances per category. For example, a 5-way-10-shot setup consists of five classes, each with ten training samples, totaling 50 training instances. We use a modified version of the ModelNet-40 [34] dataset for this task, where classes and training instances are sampled from its 40 categories.
@@
PointNet [20] - 80.39 83.7
PointNet++ [21] - 81.85 85.1
DGCNN [30] - 82.33 85.2
APES [31] - 83.67 85.8
@@
OcCo [29] 27.09 83.42 85.1
MaskPoint [17] - 85.60 86.0
ACT [5] 27.06 84.66 86.1
Point-BERT [37] 27.09 84.11 85.6
Point-MAE [19] 27.06 84.19 86.1
@@
IDPT [38] 5.69 83.50 85.3
DAPT [44] 5.65 83.83 85.5
@@
IDPT [38] 5.69 83.79 85.7
DAPT [44] 5.65 84.01 85.7
@@
Table 3: Part segmentation on the ShapeNetPart [36]. The mIoU for all classes (Cls.) and instances (Inst.) are reported. #T-Params represents the trainable parameters.
We report the ShapeNetPart [36] segmentation performance of all of the fine-tuning approaches with mIoU of all classes and instances as in Table 3. Our method has 3.45 million parameters which is a $\sim$40% reduction in the parameter size of the baseline methods, establishing itself as the most efficient method. Despite the smaller model size, GFT competes with the existing methods.
@@ 7 Conclusion and Limitation
Our findings also indicate the need for a better pretraining dataset than ShapeNet [7], which consists of synthetic data. While our method, along with other FFTs and PEFTs, performs well on downstream tasks involving synthetic datasets like ModelNet40 [34], its performance is less pronounced with real-world object scans from datasets such as ScanObjectNN [27]. This suggests that improving the pretraining process with a more diverse and representative distribution of real-world scans could further enhance the robustness and generalizability of these methods
@@ 8 Additional Ablation Study
We explore combinations of three options for projecting features from Point Transformers for classification: $T_{cls}$, patch pooling, and prompt pooling. Previous fine-tuning methods primarily relied on $T_{cls}$ and patch pooling as projections for the task head [19, 44, 38, 5, 37]. However, GFT has learnable prompts where task-specific features are accumulated. We achieve the best results by leveraging all three projection options, as shown in Table 10.
@@ 9 Training Details
Table 11 provides the comprehensive training details for our method across different classification and segmentation tasks. Classification models have been optimized using cross-entropy loss, while segmentation models utilize negative log-likelihood loss. We follow a similar training approach to earlier PEFT methods [44, 38].

### v2

@@ 4.2 Classification
For fair and robust comparison, we reproduce the results from IDPT [zha2023instance] and DAPT [zhou2024dynamic] under the same experimental setting as ours. In Table 1, we report the average overall accuracy (OA%) of models with five consecutive seeds within $[0,4]$. We design the task head to take inputs from CLS token $T_{cls}$ and pooled prompt and patch tokens from the last layer of the transformer. Two datasets are used for evaluations: ScanObjectNN for real-world object classification and ModelNet-40 for synthetic object classification.
ScanObjectNN [uy2019revisiting] is a real-world point cloud dataset containing scanned objects from indoor scenes. It consists of approximately 15K object scans with background clutter and occlusions, categorized into 15 object types. We use three dataset splits: OBJ_BG, OBJ_ONLY, and PB_T50_RS. For preprocessing, we sample 2K points per instance and apply a simple rotation, inspired by ACT [dong2023autoencoders].
ModelNet-40 [wu20153d] consists of 12,311 CAD models of synthetic objects across 40 different categories. The point clouds in these models are clean, noise-free, and uniformly distributed with independent and identical distributions. Following the standard approach, we split the dataset into 9,843 instances for training and 2,468 for testing. During training, we sample 1K points from the distribution and apply random scaling and translation.
@@ 4.3 Segmentation
For segmentation, we design a transformer decoder with skip connections from the 3rd, 6th, 9th, and 11th layers of the pretrained model. Architecture used for the decoder of the segmentation is provide in the supplementary (Section 10). In line with the baseline papers IDPT [zha2023instance] and DAPT [zhou2024dynamic], we report the best performance across ten consecutive seeds, $[0,9]$, as reproducing the baselines to compute mean metrics was computationally demanding. The evaluation dataset for the segmentation task is ShapeNetPart.
ShapeNetPart [yi2016scalable] is a part segmentation dataset containing 50 different parts across 15 object categories. Following the standard benchmarking protocol [zhou2024dynamic, zha2023instance], we divide the dataset into training and test splits, consisting of 14K and 2.9K scans, respectively. Each data instance is represented by 2K sampled points.
@@ 5 Results
The baseline foundational models favor clean and uniform point cloud data, similar to the pretraining dataset ShapeNet [chang2015shapenet]. However, real-world point clouds have a different distribution due to noise, missing data, and non-uniform point density. Thus, we benchmark with ScanObjectNN [uy2019revisiting].
As shown in Table 1, our method achieves an average performance gain of 4.12% over the FFT baseline of Point-BERT and 2.47% over Point-MAE. However, compared to the ACT baseline, there is an average performance drop of 0.5%. This could be attributed to ACT’s multi-stage pretraining [dong2023autoencoders], which, while effective, is computationally inefficient.
@@
ModelNet40 [wu20153d] contains synthetic models of objects. As illustrated in Table 1, our method does not exceed the FFT baselines of the pretrained backbones but remains competitive, with only a 0.5% average accuracy deficit. This finding of FFT can be attributed to the high similarity between the pretraining and fine-tuning datasets, as both consist of synthetic CAD models with a clean and uniform point distribution. Notably, our method outperforms other PEFT approaches across all three pretrained backbones.
@@
DGCNN-OcCo [wang2021unsupervised] 90.6 $\pm$ 2.8 92.5 $\pm$ 1.9 82.9 $\pm$ 1.3 86.5 $\pm$ 2.2
Transformer-OcCo [yu2022point] 94.0 $\pm$ 3.6 95.9 $\pm$ 2.3 89.4 $\pm$ 5.1 92.4 $\pm$ 4.6
MaskPoint [liu2022masked] 95.0 $\pm$ 3.7 97.2 $\pm$ 1.7 91.4 $\pm$ 4.0 93.4 $\pm$ 3.5
EPCL[huang2022frozen] 95.1 $\pm$ 2.7 97.3 $\pm$ 1.6 91.1 $\pm$ 4.2 93.5 $\pm$ 3.8
Point-M2AE [zhang2022point] 96.8 $\pm$ 1.8 98.3 $\pm$ 1.4 92.3 $\pm$ 4.5 95.0 $\pm$ 3.0
Point-MAE [pang2022masked] 96.3 $\pm$ 2.5 97.8 $\pm$ 1.8 92.6 $\pm$ 4.1 95.0 $\pm$ 3.0
Point-BERT [yu2022point] 94.6 $\pm$ 3.1 96.3 $\pm$ 2.7 91.0 $\pm$ 5.4 92.7 $\pm$ 5.1
ACT [dong2023autoencoders] 96.8 $\pm$ 2.3 98.0 $\pm$ 1.4 93.3 $\pm$ 4.0 95.6 $\pm$ 2.8
@@
IDPT [zha2023instance] 97.3 $\pm$ 2.1 97.9 $\pm$ 1.1 92.8 $\pm$ 4.1 95.4 $\pm$ 2.9
DAPT [zhou2024dynamic] 96.8 $\pm$ 1.8 98.0 $\pm$ 1.0 93.0 $\pm$ 3.5 95.5 $\pm$ 3.2
@@
IDPT [zha2023instance] 96.0 $\pm$ 1.7 97.2 $\pm$ 2.6 91.9 $\pm$ 4.4 93.6 $\pm$ 3.5
DAPT [zhou2024dynamic] 95.8 $\pm$ 2.1 97.3 $\pm$ 1.3 92.2 $\pm$ 4.3 94.2 $\pm$ 3.4
@@
IDPT [zha2023instance] 96.4 $\pm$ 2.5 98.3 $\pm$ 1.6 92.5 $\pm$ 4.5 95.6 $\pm$ 2.9
DAPT [zhou2024dynamic] 96.0 $\pm$ 3.0 97.4 $\pm$ 2.2 90.5 $\pm$ 4.9 94.2 $\pm$ 3.6
@@
Table 2: Few-shot learning on ModelNet40 [wu20153d]. We report the average overall accuracy (%) with the standard deviation (%) of 10 independent experiments.
Following previous works [zhou2024dynamic, zha2023instance, dong2023autoencoders, pang2022masked, yu2022point], we evaluate the pretrained model on N-way-K-shot few-shot classification. In this setting, $N\in\{5,10\}$ represents the number of object categories, while $K\in\{10,20\}$ denotes the number of training instances per category. For example, a 5-way-10-shot setup consists of five classes, each with ten training samples, totaling 50 training instances. We use a modified version of the ModelNet-40 [wu20153d] dataset for this task, where classes and training instances are sampled from its 40 categories.
@@
PointNet [qi2017pointnet] - 80.39 83.7
PointNet++ [qi2017pointnet++] - 81.85 85.1
DGCNN [wang2019dynamic] - 82.33 85.2
APES [wu2023attention] - 83.67 85.8
@@
OcCo [wang2021unsupervised] 27.09 83.42 85.1
MaskPoint [liu2022masked] - 85.60 86.0
ACT [dong2023autoencoders] 27.06 84.66 86.1
Point-BERT [yu2022point] 27.09 84.11 85.6
Point-MAE [pang2022masked] 27.06 84.19 86.1
@@
IDPT [zha2023instance] 5.69 83.50 85.3
DAPT [zhou2024dynamic] 5.65 83.83 85.5
@@
IDPT [zha2023instance] 5.69 83.79 85.7
DAPT [zhou2024dynamic] 5.65 84.01 85.7
@@
Table 3: Part segmentation on the ShapeNetPart [yi2016scalable]. The mIoU for all classes (Cls.) and instances (Inst.) are reported. #T-Params represents the trainable parameters.
We report the ShapeNetPart [yi2016scalable] segmentation performance of all of the fine-tuning approaches with mIoU of all classes and instances as in Table 3. Our method has 3.45 million parameters which is a $\sim$40% reduction in the parameter size of the baseline methods, establishing itself as the most efficient method. Despite the smaller model size, GFT competes with the existing methods.
@@ 7 Conclusion and Limitation
Our findings also indicate the need for a better pretraining dataset than ShapeNet [girshick2014rich], which consists of synthetic data. While our method, along with other FFTs and PEFTs, performs well on downstream tasks involving synthetic datasets like ModelNet40 [wu20153d], its performance is less pronounced with real-world object scans from datasets such as ScanObjectNN [uy2019revisiting]. This suggests that improving the pretraining process with a more diverse and representative distribution of real-world scans could further enhance the robustness and generalizability of these methods
@@ 8 Additional Ablation Study
We explore combinations of three options for projecting features from Point Transformers for classification: $T_{cls}$, patch pooling, and prompt pooling. Previous fine-tuning methods primarily relied on $T_{cls}$ and patch pooling as projections for the task head [pang2022masked, zhou2024dynamic, zha2023instance, dong2023autoencoders, yu2022point]. However, GFT has learnable prompts where task-specific features are accumulated. We achieve the best results by leveraging all three projection options, as shown in Table 10.
@@ 9 Training Details
Table 11 provides the comprehensive training details for our method across different classification and segmentation tasks. Classification models have been optimized using cross-entropy loss, while segmentation models utilize negative log-likelihood loss. We follow a similar training approach to earlier PEFT methods [zhou2024dynamic, zha2023instance].